OpenStudy (amistre64):

Why is: d/dx (x) = 1 :not an application of the power rule of derivatives?

myininaya (myininaya):

the slope of y=x is 1duh

OpenStudy (amistre64):

nice answer; but doesnt quite fit the question :)

myininaya (myininaya):

i don't understand the question then

myininaya (myininaya):

teach me amistre

OpenStudy (anonymous):

Uh.. I'm not sure how it's not an application of the power rule... \[\frac{d}{dx} x = \frac{d}{dx}x^1 = 1x^{1-1} = 1x^0 = 1\]

OpenStudy (amistre64):

i asked me new calc teacher today why we have:\[Dx(x)=1\] and \[Dx(x^n)=n*x^{n-1}\] as 2 seperate rules

OpenStudy (zarkon):

0^0

OpenStudy (amistre64):

she said it had something to do with the possibility of creating an indeterminate form

OpenStudy (anonymous):

Ah

OpenStudy (amistre64):

that was it

myininaya (myininaya):

zarkon is so smart

OpenStudy (amistre64):

i still dont see how that applies tho; care to provide some erudition ?

OpenStudy (anonymous):

Because if that was how it was done you'd have the derivative not existing at x=0 for the line y=x

OpenStudy (amistre64):

ohhh.... so its applicable to x=0, makes more sense :)

OpenStudy (anonymous):

Because \[x^0 = 1 \forall x \ne 0\] But at x=0 we can't evaluate it.

myininaya (myininaya):

i know 0^0 is indeterminate but why do some say =1

OpenStudy (amistre64):

so we define it to simply get rid of the "x" and save ourselves the headache :)

OpenStudy (zarkon):

use the definition of the derivative

myininaya (myininaya):

did you enjoy dance amistre?

OpenStudy (zarkon):

\[\frac{(0+h)-(0)}{h}=1\]

OpenStudy (amistre64):

lol .... the dance was ok :)

myininaya (myininaya):

ok i have a question of ethics and possibly profit...

OpenStudy (amistre64):

express it in mathematical notation and ill accept it :)

myininaya (myininaya):

So I bought a book from amazon and it cost me 92 dollars and they send me the wrong book. It costs like 172 dollars (I looked it up). Should I sell this book to possibly make a 80 dollar profit or should I make arrangements with the seller who sold me the book so they can get their book back and I can get my 92 dollars back.

myininaya (myininaya):

oops ?

OpenStudy (anonymous):

line breaks are broken.

OpenStudy (amistre64):

yes they are\[\] but im a genius lol

myininaya (myininaya):

the lines are broken?

OpenStudy (amistre64):

if they say you can keep the book, do what you want afterwards

OpenStudy (amistre64):

test test test these are on different lines

OpenStudy (amistre64):

see\[\] carriage returns have been down all day

OpenStudy (amistre64):

if they sent you the wrong book, then its not rightfully yours; try to correct the error and if they say keep it; your home free

OpenStudy (amistre64):

if p then q; if not q, then not p lol

myininaya (myininaya):

i don't really want the book

myininaya (myininaya):

i want the book i paid for lol